# Information-efficient decoding of surface codes

### v1

@@ 1 Introduction
Surface codes [kitaev97, kitaev03, dennis02] are the most well studied among error-correcting codes for quantum computing. That nearest-neighbor interactions suffice, coupled with the not-too-stringent noise threshold, have made the planar variant [bravyi98, fowler12] the code of choice for many hardware architectures. The past three years have seen experimental demonstrations of beyond-breakeven error correction using surface codes [google2023suppressing, eickbusch2025demonstration, google2025quantum, bluvstein2024logical, bluvstein2025architectural]. The success of surface codes also motivated the current intense search for good quantum low-density parity-check (q-LDPC) codes with similar locality and threshold properties but with better coding rates [breuckmann2021, bravyi2024high, xu2024constant, pattison2025hierarchical, berthusen2025toward].
@@
Error correction with surface codes, as is true with stabilizer codes in general, demands repeated measurements of the stabilizer generators that define the code. These measurements generate syndrome bits across the entire code lattice, and these bits are processed by a classical decoder for error diagnosis. Many decoding algorithms exist, including the traditional minimum-weight perfect-matching (MWPM) via the Blossom algorithm [edmonds1965] (implemented in the popular PyMatching package [higgott22]), the union-find approach [delfosse2021], the Sparse-Blossom variant of MWPM [higgott25], the maximum-likelihood decoder [bravyi2014], as well as machine-learning decoding protocols (e.g., Refs. [torlai2017, krastanov2017]); see also Ref. [iOlius2024] for a review on surface-code decoders. Each of these decoding algorithms takes in the measured syndrome bits and produces a best guess for the errors that have occurred to generate those syndromes.
@@
For use within a quantum computation, the decoders must be able to keep up with the syndrome generation, to not suffer from the so-called backlog problem [terhal2015quantum] that can lead to exponential slowdown of the computation, negating any quantum speed advantage one may hope to achieve. Syndromes have to be processed as they are generated [terhal2015quantum, skoric2023parallel, tan2023scalable], to deduce the correct Pauli frame needed at the moment of application of a logical $T$ gate (via the magic-state approach [bravyi2005]) during the computation. This puts a requirement on the classical communication rate between the quantum processing unit and the classical computer that performs the decoding.
@@
Past works [das_afs_2022, zhang_classical_2023, camps_evaluation_2024] have analyzed the communication bandwidth requirements. In every syndrome measurement cycle, for each logical qubit encoded with distance-$d$ rotated planar surface code, $O(d^{2})$ bits of syndrome information are generated—the number of syndrome bits scales as the area of the $d\times d$ lattice. The transmission of the $d^{2}$ bits of information per logical qubit has to be completed within the syndrome measurement cycle time to keep up with the syndrome generation rate. This translates into gigabits or even terabits per second of classical data transmission for typical noise and reasonably sized quantum algorithms [das_afs_2022]; for comparison, a typical server stack can handle perhaps terabits per second of communication. Handling such large data transmission rates requires multiple high-bandwidth classical communication channels, but too many such connections between the (typically) cryogenic quantum processing unit (QPU) where the syndromes are generated and the hot classical processing unit (CPU) outside will ultimately degrade the quality of the QPU.
@@
Ref. [das_afs_2022] proposed data compression protocols to reduce this requisite data transmission rate, noting that nontrivial syndromes arise at a rate that scales with the error probability. Yet, any such data compression, to actually reduce the data transmission rate, has to be done within the QPU, hence limiting the protocols to only those simple enough for the on-chip CPU (perhaps an FPGA with only basic Boolean operations) to handle. Furthermore, one cannot guarantee efficient compression in all cases since the syndrome data can, in principle, be completely random.
@@
We thus ask the question, before we attempt such data compression, if all the syndrome data are necessary for decoding in the first place. Can we have more information-efficient decoders instead? Here, we answer in the affirmative for planar surface codes, and introduce two information-efficient decoders: the row-column decoder and the boundary decoder. The row-column decoder aggregates the syndrome information in a row or column—simple enough for an on-chip FPGA to handle—and only those aggregates are transmitted to and processed by the classical decoding algorithm (e.g., MWPM). The boundary decoder relies on the dynamic syndrome measurement circuits of Ref. [mcewen23] to “pump" errors to the boundaries of the surface-code lattice such that only the syndrome bits at the boundaries have to be extracted to the decoding algorithm for error diagnosis. In both cases, the number of transmitted syndrome bits scales with $d$, the width of the surface-code lattice, rather than with its area $d^{2}$.
@@ 3.1 Row-column decoder
The goal of the decoder is thus to enforce an even parity for $X(Z)$ errors along each row(column) in the surface-code patch, by detecting those rows(columns) with an odd number of $X(Z)$ errors, corresponding to an odd number of defects in the neighboring ancillary-qubit rows(columns). We thus aggregate the measurement outcomes of all the $Z$ stabilizers in the same row, or all the $X$ stabilizers in the same column, into a single “detecting region" [mcewen23]. The aggregated row/column outcome (or detector) corresponds to taking the parity of the measurement outcomes of the individual $Z(X)$ stabilizers within a specific row(column) in the current and preceding syndrome extraction rounds—the temporal comparison identifies the defects, and taking the parity thereafter identifies when there is an odd number of defects. The number of such detecting regions is 2(d1)2(d-1) per syndrome measurement round, with d1d-1 of them dedicated to detecting (odd number of) $X$ errors and d1d-1 for $Z$ errors. From the aggregated parity values, standard MWPM decoding is used to pair up those parity values, and diagnosed errors can be attributed to data qubits on the boundary. This would finally give an even number of $X$ and $Z$ errors in every row and column, respectively.
@@ 3.2 Boundary decoder
The row-column decoder uses syndrome information gathered from the standard SE circuits of Fig. 1(c). For our second information-efficient decoder, we rely instead on the dynamic SE circuits introduced in Ref. [mcewen23]. Dynamic SE circuits give rise to novel features in the error correction process compared to traditional circuits. In traditional SE circuits, stabilizers remain static and in place during the syndrome measurement rounds; in dynamic circuits, the stabilizers evolve by expanding and contracting, or even moving around. This evolution allows the flexibility of changing the code generators after each SE round, and Ref. [mcewen23] made use of this flexibility to reduce the connectivity requirements for implementing SE for the rotated planar surface codes. Recent experimental work has demonstrated the feasibility of this circuit in hardware [eickbusch2025demonstration]. Dynamic circuits have further been applied to lower connectivity demands in bivariate bicycle codes [shaw_lowering_2025], and to provide resilience against qubit dropouts [debroy2024luci].
A curious side effect of the dynamic SE circuits is that errors present before SE can propagate into different errors afterwards. For the SE to not break the requirements of fault tolerance and affect the error-correction performance, dynamic SE circuits should not spread errors uncontrollably. However, they are not constrained to keep the errors in place. In fact, the circuit of Ref. [mcewen23] has the effect of moving errors along the benign directions towards the boundaries.
@@
Figure 2: The 3CX (or hex-grid) syndrome extraction circuit of Ref. [mcewen23]. (a) The syndrome extraction circuit, comprising two halves, A (left half) and B (right half), applied in alternation to each square in the bulk of the surface-code patch. The purple square on the right gives the data qubit labeling; the ancilla is the black dot in the middle. (b) and (c) show how input errors on the data qubits are moved to different qubits after the A- and B-circuits. The colored $X$s or $Z$s indicate where the error has spread to after the preceding gate operation. Only input errors that are moved in the current SE round are shown. The A- and B-circuits pump errors along the benign directions towards the boundaries of the surface-code patch; see main text for more details.
@@
To see this, we look more closely at the dynamic SE circuit proposed in Ref. [mcewen23], shown here in Fig. 2(a). This is the hex-grid circuit of Ref. [mcewen23], referred to elsewhere also as the “3CX” circuit [hetenyi24], since it requires CNOT (or CX) connections between three—not the usual four—data qubits (labeled 1–3 in Fig. 2) and the ancilla. For each square in the bulk of the surface-code patch comprising four data qubits at its corners (purple square in top right of Fig. 2), the dynamic SE circuit alternates between the A-circuit [left half of Fig. 2(a)] and the B-circuit (right half). At initial time, if the square corresponds to a $X(Z)$ stabilizer [red(blue) square in Fig. 1], it begins with the A(B)-circuit; the A- or B-circuits are subsequently applied alternately on those squares.
@@
Viewed in terms of how the code stabilizers morph as SE proceeds, the A(B)-circuit changes a $X(Z)$-stabilizer [red(blue)] square into a $Z(X)$-stabilizer [blue(red)] square. Each square in the checkerboard pattern of the code patch thus alternates between being an $X$ stabilizer and a $Z$ stabilizer, switching every half cycle of the full SE round (A+B or B+A). The A(B)-circuit is applied whenever the square is an $X(Z)$ stabilizer. Every $X$ or $Z$ stabilizer in the code patch is measured in each round, either in A or B. This structure becomes clearer when analyzed via detecting regions. A detector in a circuit is a set of measurement outcomes with deterministic parity in the absence of errors. The detecting region associated with that detector is the set of circuit locations such that Pauli errors anywhere in the set would flip that parity value. At the beginning of an SE round (A or B), there are $2(d^{2}-1)$ detecting regions in all, with $d^{2}-1$ expanding and $d^{2}-1$ contracting. The expanding detecting regions originate from ancilla resets and evolve to cover the code stabilizers by the end of the SE round. Conversely, the contracting detecting regions are those that began expanding in the previous round and terminate in ancilla measurements in the current round. At the start of a round, the $d^{2}-1$ contracting regions collectively cover the $d^{2}-1$ stabilizer generators of the code. By the end of the round, the measurements of the ancillary qubits within those regions yield the eigenvalues of the stabilizers. See Ref. [mcewen23] for more details.
@@
From Figs. 2(b) and (c), we see that the dynamic SE circuits move the errors. Specifically, they “pump” $X$ errors rightwards—along the benign-$X$ direction—by one unit (e.g., qubit 1 to 2) every SE round; at the same time, $Z$ errors are pumped downwards—along the benign-$Z$ direction—by one unit (e.g., qubit 1 to 3). Eventually, an error originating anywhere in surface-code patch reaches the right or bottom boundaries of the patch. At the boundary, the stabilizers and SE circuits for the boundary data qubits have to be chosen carefully, as highlighted in Ref. [mcewen23]; we leave the reader to consult the original reference for the full details. Of relevance for us here is the fact that the specific choice in Ref. [mcewen23] causes errors that arrive at the right and bottom boundaries to remain there, unchanged by subsequent SE rounds, except when an identical error arrives at the same location and neutralizes an existing error there. These error movements are visualized in a video available online [my_2025_17851272].
@@ 4 Decoder performance
The memory experiments were simulated using Stim [gidney21], with PyMatching [higgott22] (with Sparse Blossom [higgott25]) as the decoding algorithm for the collected syndrome information. For the row-column decoder, only the aggregated row/column parity values were sent to the decoder; detectors were formed by comparing the current round’s parity values with those from the same row/column in the previous SE round. For the boundary decoder, only the syndrome bits on the right and bottom boundaries were sent to the decoder; detectors were formed by comparing the current syndrome outcome at a boundary location with the syndrome value at the same location in the previous round.
@@ 5 Conclusion
We have described two decoders that use only a reduced set of the spatial syndrome information. Despite the higher logical error probability compared to the areal decoder, our information-efficient decoders give a significant reduction in the amount of information that has to be transmitted out to the classical decoding algorithm, from $O(d^{2})$ per SE round for the areal decoder to $O(d)$ for the information-efficient decoders, alleviating the demands on the classical communication rate needed to extract the syndrome information to the decoder on the CPU. Furthermore, the size of the decoding graph used in PyMatching is also correspondingly reduced from $O(d^{2})$ to $O(d)$, leading to a significant speedup in the decoding runtime, from $O(d^{6}\log(d))$ to $O(d^{3}\log(d))$ [higgott22].

### v2

@@ 1 Introduction
Surface codes [24, 25, 16] are the most well studied among error-correcting codes for quantum computing. That nearest-neighbor interactions suffice, coupled with the not-too-stringent noise threshold, have made the planar variant [6, 19] the code of choice for many hardware architectures. The past three years have seen experimental demonstrations of beyond-breakeven error correction using surface codes [2, 18, 1, 4, 5]. The success of surface codes also motivated the current intense search for good quantum low-density parity-check (q-LDPC) codes with similar locality and threshold properties but with better coding rates [10, 7, 35, 29, 3].
@@
Error correction with surface codes, as is true with stabilizer codes in general, demands repeated measurements of the stabilizer generators that define the code. These measurements generate syndrome bits across the entire code lattice, and these bits are processed by a classical decoder for error diagnosis. Many decoding algorithms exist, including the traditional minimum-weight perfect-matching (MWPM) via the Blossom algorithm [17] (implemented in the popular PyMatching package [23]), the union-find approach [14], the Sparse-Blossom variant of MWPM [22], the maximum-likelihood decoder [9], as well as machine-learning decoding protocols (e.g., Refs. [34, 26]); see also Ref. [15] for a review on surface-code decoders. Each of these decoding algorithms takes in the measured syndrome bits and produces a best guess for the errors that have occurred to generate those syndromes.
@@
For use within a quantum computation, the decoders must be able to keep up with the syndrome generation, to not suffer from the so-called backlog problem [33] that can lead to exponential slowdown of the computation, negating any quantum speed advantage one may hope to achieve. Syndromes have to be processed as they are generated [33, 31, 32], to deduce the correct Pauli frame needed at the moment of application of a logical $T$ gate (via the magic-state approach [8]) during the computation. This puts a requirement on the classical communication rate between the quantum processing unit and the classical computer that performs the decoding.
@@
Past works [12, 36, 11] have analyzed the communication bandwidth requirements. In every syndrome measurement cycle, for each logical qubit encoded with distance-$d$ rotated planar surface code, $O(d^{2})$ bits of syndrome information are generated—the number of syndrome bits scales as the area of the $d\times d$ lattice. The transmission of the $d^{2}$ bits of information per logical qubit has to be completed within the syndrome measurement cycle time to keep up with the syndrome generation rate. This translates into gigabits or even terabits per second of classical data transmission for typical noise and reasonably sized quantum algorithms [12]; for comparison, a typical server stack can handle perhaps terabits per second of communication. Handling such large data transmission rates requires multiple high-bandwidth classical communication channels, but too many such connections between the (typically) cryogenic quantum processing unit (QPU) where the syndromes are generated and the hot classical processing unit (CPU) outside will ultimately degrade the quality of the QPU.
@@
Ref. [12] proposed data compression protocols to reduce this requisite data transmission rate, noting that nontrivial syndromes arise at a rate that scales with the error probability. Yet, any such data compression, to actually reduce the data transmission rate, has to be done within the QPU, hence limiting the protocols to only those simple enough for the on-chip CPU (perhaps an FPGA with only basic Boolean operations) to handle. Furthermore, one cannot guarantee efficient compression in all cases since the syndrome data can, in principle, be completely random.
@@
We thus ask the question, before we attempt such data compression, if all the syndrome data are necessary for decoding in the first place. Can we have more information-efficient decoders instead? Here, we answer in the affirmative for planar surface codes, and introduce two information-efficient decoders: the row-column decoder and the boundary decoder. The row-column decoder aggregates the syndrome information in a row or column—simple enough for an on-chip FPGA to handle—and only those aggregates are transmitted to and processed by the classical decoding algorithm (e.g., MWPM). The boundary decoder relies on the dynamic syndrome measurement circuits of Ref. [27] to “pump" errors to the boundaries of the surface-code lattice such that only the syndrome bits at the boundaries have to be extracted to the decoding algorithm for error diagnosis. In both cases, the number of transmitted syndrome bits scales with $d$, the width of the surface-code lattice, rather than with its area $d^{2}$.
@@ 3.1 Row-column decoder
The goal of the decoder is thus to enforce an even parity for $X(Z)$ errors along each row(column) in the surface-code patch, by detecting those rows(columns) with an odd number of $X(Z)$ errors, corresponding to an odd number of defects in the neighboring ancillary-qubit rows(columns). We thus aggregate the measurement outcomes of all the $Z$ stabilizers in the same row, or all the $X$ stabilizers in the same column, into a single “detecting region" [27]. The aggregated row/column outcome (or detector) corresponds to taking the parity of the measurement outcomes of the individual $Z(X)$ stabilizers within a specific row(column) in the current and preceding syndrome extraction rounds—the temporal comparison identifies the defects, and taking the parity thereafter identifies when there is an odd number of defects. The number of such detecting regions is 2(d1)2(d-1) per syndrome measurement round, with d1d-1 of them dedicated to detecting (odd number of) $X$ errors and d1d-1 for $Z$ errors. From the aggregated parity values, standard MWPM decoding is used to pair up those parity values, and diagnosed errors can be attributed to data qubits on the boundary. This would finally give an even number of $X$ and $Z$ errors in every row and column, respectively.
@@ 3.2 Boundary decoder
The row-column decoder uses syndrome information gathered from the standard SE circuits of Fig. 1(c). For our second information-efficient decoder, we rely instead on the dynamic SE circuits introduced in Ref. [27]. Dynamic SE circuits give rise to novel features in the error correction process compared to traditional circuits. In traditional SE circuits, stabilizers remain static and in place during the syndrome measurement rounds; in dynamic circuits, the stabilizers evolve by expanding and contracting, or even moving around. This evolution allows the flexibility of changing the code generators after each SE round, and Ref. [27] made use of this flexibility to reduce the connectivity requirements for implementing SE for the rotated planar surface codes. Recent experimental work has demonstrated the feasibility of this circuit in hardware [18]. Dynamic circuits have further been applied to lower connectivity demands in bivariate bicycle codes [30], and to provide resilience against qubit dropouts [13].
@@
Figure 2: The 3CX (or hex-grid) syndrome extraction circuit of Ref. [27]. (a) The syndrome extraction circuit, comprising two halves, A (left half) and B (right half), applied in alternation to each square in the bulk of the surface-code patch. The purple square on the right gives the data qubit labeling; the ancilla is the black dot in the middle. (b) and (c) show how input errors on the data qubits are moved to different qubits after the A- and B-circuits. The colored $X$s or $Z$s indicate where the error has spread to after the preceding gate operation. Only input errors that are moved in the current SE round are shown. The A- and B-circuits pump errors along the benign directions towards the boundaries of the surface-code patch; see main text for more details. (d) shows an example in the 2D layout for 1 SE round, with a $Y$ error (purple circle) occurring in the bulk just before the start of the round; the two panels show how the $X$ (red circle) and $Z$ (blue circle) errors propagate under the two halves of the SE round under consecutive AA and $B$ circuits. The $A/B$ circuit responsible for the movement is indicated by the arrow beside the $A/B$ labels.
A curious side effect of the dynamic SE circuits is that errors present before SE can propagate into different errors afterwards. For the SE to not break the requirements of fault tolerance and affect the error-correction performance, dynamic SE circuits should not spread errors uncontrollably. However, they are not constrained to keep the errors in place. In fact, the circuit of Ref. [27] has the effect of moving errors along the benign directions towards the boundaries.
@@
To see this, we look more closely at the dynamic SE circuit proposed in Ref. [27], shown here in Fig. 2(a). This is the hex-grid circuit of Ref. [27], referred to elsewhere also as the “3CX” circuit [21], since it requires CNOT (or CX) connections between three—not the usual four—data qubits (labeled 1–3 in Fig. 2) and the ancilla. For each square in the bulk of the surface-code patch comprising four data qubits at its corners (purple square in top right of Fig. 2), the dynamic SE circuit alternates between the A-circuit [left half of Fig. 2(a)] and the B-circuit (right half). At initial time, if the square corresponds to a $X(Z)$ stabilizer [red(blue) square in Fig. 1], it begins with the A(B)-circuit; the A- or B-circuits are subsequently applied alternately on those squares.
@@
Viewed in terms of how the code stabilizers morph as SE proceeds, the A(B)-circuit changes a $X(Z)$-stabilizer [red(blue)] square into a $Z(X)$-stabilizer [blue(red)] square. Each square in the checkerboard pattern of the code patch thus alternates between being an $X$ stabilizer and a $Z$ stabilizer, switching every half cycle of the full SE round (A+B or B+A). The A(B)-circuit is applied whenever the square is an $X(Z)$ stabilizer. Every $X$ or $Z$ stabilizer in the code patch is measured in each round, either in A or B. This structure becomes clearer when analyzed via detecting regions. A detector in a circuit is a set of measurement outcomes with deterministic parity in the absence of errors. The detecting region associated with that detector is the set of circuit locations such that Pauli errors anywhere in the set would flip that parity value. At the beginning of an SE round (A or B), there are $2(d^{2}-1)$ detecting regions in all, with $d^{2}-1$ expanding and $d^{2}-1$ contracting. The expanding detecting regions originate from ancilla resets and evolve to cover the code stabilizers by the end of the SE round. Conversely, the contracting detecting regions are those that began expanding in the previous round and terminate in ancilla measurements in the current round. At the start of a round, the $d^{2}-1$ contracting regions collectively cover the $d^{2}-1$ stabilizer generators of the code. By the end of the round, the measurements of the ancillary qubits within those regions yield the eigenvalues of the stabilizers. See Ref. [27] for more details.
@@
From Figs. 2(b) and (c), we see that the dynamic SE circuits move the errors. Specifically, they “pump” $X$ errors rightwards—along the benign-$X$ direction—by one unit (e.g., qubit 1 to 2) every SE round; at the same time, $Z$ errors are pumped downwards—along the benign-$Z$ direction—by one unit (e.g., qubit 1 to 3). Eventually, an error originating anywhere in surface-code patch reaches the right or bottom boundaries of the patch. At the boundary, the stabilizers and SE circuits for the boundary data qubits have to be chosen carefully, as highlighted in Ref. [27]; we leave the reader to consult the original reference for the full details. Of relevance for us here is the fact that the specific choice in Ref. [27] causes errors that arrive at the right and bottom boundaries to remain there, unchanged by subsequent SE rounds, except when an identical error arrives at the same location and neutralizes an existing error there. These error movements are visualized in a video available online [28].
@@ 4 Decoder performance
The memory experiments were simulated using Stim [20], with PyMatching [23] (with Sparse Blossom [22]) as the decoding algorithm for the collected syndrome information. For the row-column decoder, only the aggregated row/column parity values were sent to the decoder; detectors were formed by comparing the current round’s parity values with those from the same row/column in the previous SE round. For the boundary decoder, only the syndrome bits on the right and bottom boundaries were sent to the decoder; detectors were formed by comparing the current syndrome outcome at a boundary location with the syndrome value at the same location in the previous round.
@@ 5 Conclusion
We have described two decoders that use only a reduced set of the spatial syndrome information. Despite the higher logical error probability compared to the areal decoder, our information-efficient decoders give a significant reduction in the amount of information that has to be transmitted out to the classical decoding algorithm, from $O(d^{2})$ per SE round for the areal decoder to $O(d)$ for the information-efficient decoders, alleviating the demands on the classical communication rate needed to extract the syndrome information to the decoder on the CPU. Furthermore, the size of the decoding graph used in PyMatching is also correspondingly reduced from $O(d^{2})$ to $O(d)$, leading to a significant speedup in the decoding runtime, from $O(d^{6}\log(d))$ to $O(d^{3}\log(d))$ [23].